OpenStudy (anonymous):

Simplify and solve 3i - 10 / 8i -5

OpenStudy (anonymous):

Is this it? \[\frac{3i-10}{8i-5}\]

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

I'm actually lost in here too. lol

OpenStudy (anonymous):

Maybe you are missing some information. You wrote down "solve" but we can't solve anything since this is not an equation it's only an expression.

OpenStudy (anonymous):

thats all the problem says

OpenStudy (anonymous):

Rationalize

OpenStudy (anonymous):

\[\frac{( 3i - 10) }{ (8i-5) }*\frac{ (8i + 5) }{ (8i+5)}\]